OpenStudy (anonymous):

Find the critical numbers of the function and describe the behavior of f(x) = x8(x - 3)7 at these numbers.

OpenStudy (anonymous):

\[f(x) = x^8(x - 3)^7\]\[f'(x)=8x^7(x-3)^7+7x^8(x-3)^6\] via hte product rule

OpenStudy (anonymous):

factor the derivative, so you can find the zeros easily

OpenStudy (anonymous):

And that's the first derivative. What is it I have to do to find the zeros? That has always confused me

OpenStudy (anonymous):

you have to write it in factored form first

OpenStudy (anonymous):

it is pretty clear just from looking at it that one zero is \(0\) and also one zero is \(3\) but you have to factor it both terms have a common factor of \(x^7(x-3)^6\) factor that out

OpenStudy (anonymous):

it is algebra at this point, i can show you the method if you like

OpenStudy (anonymous):

I'm not good at algebra, my strong points are in geometry, if you would please I would like that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f'(x)=8x^7(x-3)^7+7x^8(x-3)^6\] common factor in both terms is \(x^7(x-3)^6\) so first step is \[8x^7(x-3)^7+7x^8(x-3)^6=x^7(x-3)^6\left(\text{something}\right)\]

OpenStudy (anonymous):

ok to this step?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok good and the "something" which you get by thinking has to be \(8(x-3)+7x\)

OpenStudy (anonymous):

Oh okay. That makes sense now.

OpenStudy (anonymous):

\[8x^7(x-3)^7+7x^8(x-3)^6=x^7(x-3)^6\left(8(x-3)+7x\right)\]then algebra in the second part, gives \[x^7(x-3)^6(8x-24+7x)=x^7(x-3)^6(15x-24)\] and now the other zero is easy to find

OpenStudy (anonymous):

And they would be... -3 and -24?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

Calculus really makes me feel stupid.

OpenStudy (anonymous):

set each factor equal to zero and solve: \[x^7=0\iff x=0\\ x-3=0\iff x=3\\ 15x-24=0\iff 15x=24\iff x=\frac{24}{15}=\frac{8}{5}\]

OpenStudy (anonymous):

that is just because you have to use all that elementary algebra you may have forgotten that is why they taught that first

OpenStudy (anonymous):

the ideas of calc are not that hard, the mechanics are the confusing part

OpenStudy (anonymous):

Yeah, the logic doesn't seem logical to me. So then are those the critical points?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. And it's increasing until 0, decreasing until 8/5, and increasing until three?

OpenStudy (anonymous):

yes, i believe it is, although it is increasing after 3 as well, the derivative does not change sign at 3 you can check the picture here http://www.wolframalpha.com/input/?i=x^8%28x+-+3%29^7+domain+-1..4

OpenStudy (anonymous):

Okay, thank you. That helped me a lot

OpenStudy (drakedoesmc):

@satellite73 please close the question in biology about taylor swift thank you and have a nice day